OpenStudy (anonymous):

If the point(x,3√3)is on the unit circle, what is x?

OpenStudy (irishboy123):

on the unit circle, \[x ^{2} + y ^ {2} = 1\]

OpenStudy (anonymous):

there are four answers

OpenStudy (anonymous):

2/3 \[\sqrt{6}/3\] 4/9 \[\sqrt{2}/3\]

OpenStudy (yolo_king):

\[x^2+y^2 =1\] Now we have \[y=\frac{ \sqrt{3} }{ 3 }\]

OpenStudy (anonymous):

those are my four options

OpenStudy (yolo_king):

So then we plug this in \[x^2+(\frac{ \sqrt{3} }{ 3})^2=1\] Giving us \[x^2+\frac{ 3}{ 9 }= x^2+\frac{ 1 }{ 3}=1\]

OpenStudy (anonymous):

i see what your saying but that isnt one of the given options

OpenStudy (yolo_king):

Thats because that's not the answer :), Solve for x And that should be your answer.

OpenStudy (anonymous):

would it be \[\sqrt{2}/3\]?